# The Popularity Hypothesis in Software Security: A Large-Scale Replication with PHP Packages

### v1

@@ Abstract
There has been a long-standing hypothesis that a software’s popularity is related to its security or insecurity in both research and popular discourse. There are also a few empirical studies that have examined the hypothesis, either explicitly or implicitly. The present work continues with and contributes to this research with a replication-motivated large-scale analysis of software written in the PHP programming language. The dataset examined contains nearly four hundred thousand open source software packages written in PHP. According to the results based on reported security vulnerabilities, the hypothesis does holds; packages having been affected by vulnerabilities over their release histories are generally more popular than packages without having been affected by a single vulnerability. With this replication results, the paper contributes to the efforts to strengthen the empirical knowledge base in cyber and software security.
keywords:
@@
Software security, software vulnerability, replication, web applications, software ecosystems, dependencies, downloads, installs, stars, forks
1 Introduction
@@
The paper examines what can be called a popularity hypothesis in empirical software security research; an assumption that a software’s popularity affects its security or insecurity. The hypothesis has been previously examined with Java packages, the conclusion being that popularity is not a reliable indicator of software security [1]. The present paper revisits this hypothesis with a large-scale analysis of open source software packages written in the PHP programming language. Given the existing work, the paper is by definition a replication.
@@
Within the domain of replication studies, the paper can be classified as a dependent or a conceptual replication: although both the data and methods are different, the paper is designed with an explicit reference to previous work in mind [2, 3, 4]. The methodological difference can be justified on the grounds that the replicated study was based only on correlation analysis [1]. In contrast, the paper uses also regression analysis and classification algorithms, which can be argued to yield stronger empirical evidence in general. The recently discussed generalizability problems in empirical software engineering [5] justify the use of a different dataset. If the hypothesis holds, or does not hold, with two distinct datasets, there is either stronger or weaker evidence for it, respectively.
@@
The underlying replication logic focuses on the previous work’s conclusion—it is important to know whether or not the hypothesis holds with a different dataset and different methods because the hypothesis is rather fundamental in many ways. It also frequently appears in popular discourse. The primary historical example would be Microsoft Windows; not so long ago, it was commonly discussed and speculated by both laypeople and professionals that the popularity of the operating system was a definite factor behinds its historical insecurity. To quote from a famous security engineering textbook, most “malware writers targeted Windows rather than Mac or Linux through the 2000s and 2010s as there are simply more Windows machines to infect”, and, furthermore, the “model replicated itself when smartphones took over the world” [6, p. 296]. Although there is no particular reason to question these assertions, it is still worth remarking that they are not actually backed by any empirical evidence in the textbook. This point motivates a need to examine the hypothesis further empirically.
@@
Given a distinction between “hard” and “soft” theories in empirical software engineering [7], the popularity hypothesis is on the side of hard theories because it presumes a causal relation between a cause (popularity) and effect (insecurity), but the theorizing involved is still rather soft. In other words, assuming that the hypothesis holds, the underlying explanations remain open to a debate.
@@
The starting point for an explanation given by what is now known as (information or cyber) security economics would be incentives [6]. That is, cyber criminals and other adversaries have an incentive to target popular software products, packages, and projects, whether the targeting is about finding vulnerabilities from those or explicitly attacking them. Then: even though both commercial software vendors and open source software projects have an incentive to develop secure code and provide defenses, it is frequently discussed in the literature that software security and cyber security in general are not prioritized because there are no pressing incentives to do so [8]. A slightly different theorization path involves the famous so-called Linus law; that all bugs are shallow when there are many eyes involved to look at code [9]. From this alternative perspective, there are also empirical studies indicating that widely used and hence popular open source software packages often contain more reported bugs and vulnerabilities than less popular packages [10, 11, 12]. These empirical studies further reinforce the need and motivation for a replication of the popularity hypothesis study already mentioned earlier. The italics placed upon the word reported are also worth briefly explaining. The reason for the italics is that empirical observations cannot generally prove that something is secure, although these can be used to make claims about insecurity [13]. This fundamental point should be kept in mind throughout the paper; only reported security issues are observed. By implication, the popularity hypothesis in itself may not necessarily tell everything, or in some cases even anything, about software security per se.
@@
    With respect to the replicated study [1], the paper examines a substantially larger dataset and uses more sophisticated methods for its analysis.
@@
The paper’s remainder is structured into a conventional format. The opening Section 2 further frames the paper with a larger scope of related work. Then, the empirical materials and the methods for examining these are elaborated in Section 3. This section also clarifies what is meant by software popularity; in general, it is operationalized as a degree to which a package is widely adopted and actively engaged within its ecosystem. The subsequent Section 4 presents the replication results after which conclusions follow in the final Section 5.
2 Related Work
@@
The paper’s reference domain is empirical software engineering focusing on software security. The domain is large, but with some suitable framings, it is possible to narrow it down. The first obvious framing is toward empirical research. The second equally obvious framing is toward the PHP programming language. With these two framings, a few examples can be given about implicitly related work. For instance, technical debt of PHP packages for web applications has been investigated empirically [14]. Regarding web applications and websites in general, another example would be a large-scale empirical study investigating the adoption of PHP interpreter releases among website deployments [15]. A third framing can be done toward software security. The previous two examples are in the scope also with this further software security framing—technical debt is related to software quality, which is a superset for software security, while the adoption (or a lack thereof) of new PHP interpreter releases contains security risks in case many (or some important) websites do not update their interpreters for a reason or another. The reason for the risks is that also the PHP interpreter has seen many vulnerabilities over the years.
@@
A lot of empirical research has also been done to examine the security of software written in PHP more generally. For instance, the security of plugins for popular PHP-based web frameworks has been investigated [16, 12]. Regarding software vulnerabilities and their underlying weaknesses, and despite the availability of defensive solutions [17], cross-site scripting and input validation more generally have been the most typical weaknesses in PHP software [12, 18]. Though, it can be noted that this observation is hardly unique to PHP. In other words, the same weaknesses typically lead the scoreboards also with other popular interpreted programming languages used in the web domain [19]. Having said that, it should be emphasized that empirical software security research on PHP applications is not limited only on vulnerabilities and weaknesses. A good example would be a forensics investigation of PHP applications in containerized cloud environments based on empirical log analysis [20]. Furthermore, a third framing can be done toward replication research.
@@
A fourth and final clarifying framing can be done toward research on software ecosystems and particularly the security or insecurity of software distributed via them. The empirical security-oriented ecosystem research is again vast. A recurring theme is that software packages distributed in software ecosystems via language-specific package managers are generally of poor equality, often containing security issues of various kinds, including concrete vulnerabilities [23]. In addition, the ecosystems exhibit a risk of malware being uploaded, as is often done together with so-called typo-squatting [19, 24]. The underlying security risks are intensified by the heavy use of software dependencies in these ecosystems [25]. Given this background, it is understandable that there are also various ongoing funded projects to improve the ecosystems and the supply chain security of open source software in general [8]. The software ecosystem research and dependencies are also important to underline because the concept of (software) popularity is closely related to them [26, 27]. The framing toward software ecosystems is also relevant as it is also directly related to the dataset examined. On that note, the materials and methods should also be elaborated.
3 Materials and Methods
@@
3.1 Data
@@
The dataset was collected on December 2024 from Packagist [28], a repository for PHP software packaged with the Composer package manager [29]. The dataset is available online for any interested party, possibly regarding further replication work [30]. In total, $n=381,993$ packages are covered in the dataset. The dataset is composed of eight simple meta-data metrics provided on the Packagist’s website: security, installs, dependents, suggesters, stars, watchers, forks, and releases. Of these metrics, all except security and releases can be interpreted to proxy slightly different dimensions of software popularity. The outlying releases metric counts the number of releases made for a given open source PHP package, as reported on the Packagist’s website. All metrics have a continuous scale.
@@
While installation amounts, installs, provide a relatively straightforward and widely used metric for software popularity [11, 26, 27, 12], the dependents metric conveys a package’s popularity in terms of incoming edges in a dependency network between packages [26]. With this metric a package is seen as popular when many other packages depend on it. Both installation (or download) counts and dependencies have also been used to help security-related decision-making about funding applications for open source software projects [8]. Although documentation is unfortunately lacking, the suggesters metric generally proxies “soft” dependencies; for each package, Packagist recommends also further packages in addition to the “hard” dependencies, as captured by dependents, without which a package does not install or function properly.
@@
The stars, watchers, and forks metrics are explicitly tied to functionality provided by GitHub on which almost all of the actual development of the PHP packages occurs today. In general, with “stars” developers can flag packages and projects they perceive as interesting or relevant. By adding oneself as a “watcher” to a project, one can get information about changes in the project. The forks metric, in turn, refers to the number of times a given project was forked, as has been common in the open source world throughout the decades [31]. According to practitioner surveys, all three metrics have been seen as relevant for proxying popularity [32]. Finally, the dependent metric is security, which counts and records vulnerabilities reported for a package. Many (but not all) of the vulnerabilities tracked on Packagist are identified with Common Exposures and Vulnerabilities (CVEs), while the associated security advisories typically point toward GitHub. In this regard, it can be noted that no additional data was retrieved to validate the CVE-referenced vulnerability information provided on Packagist. Nor was GitHub queried to retrieve further software development data.
3.2 Methods
@@
The primary methodology is based on binary classification. In addition, a few descriptive statistics and a brief ordinary least squares (OLS) regression analysis are presented. As these do not require a particular exposition, the following discussion concentrates on the classification.
3.2.1 Classification
@@
Imbalance is a highly typical problem in computational cyber security applications [33, 34]. Although not a problem for the current paper, not only is cyber security data often imbalanced but it also often contains missing values [35]. Even without a problem of missing values, as will be soon seen in Section 4, the dataset examined can be characterized to contain extreme imbalance. In other words, only a small minority of the PHP packages examined has been affected by one or more vulnerabilities.
@@
To tackle the issue, four computational solutions are used. The first is oversampling or upsampling, as it is also known. It balances a training set by replicating the minority class, which in the present context refers to packages that have been affected by one or more vulnerabilities during their entire release histories. The second is undersampling or downsampling; it eliminates observations from the majority class (non-vulnerable packages) until a training set is balanced. Although there are variants that use synthetic data [33], both the oversampling and undersampling solutions are computed with random samples.
@@
3.2.2 Performance Metrics
Imbalance affects also metrics for evaluating classification performance. Among other things [39], the traditional accuracy metric is biased and misleading in this context. Although the $F_{1}$ metric is often used as an alternative, it too can be problematic in cyber security applications [45]. Therefore, three alternative performance metrics are used.
@@ 3.2.2 Performance Metrics
Third, the conventional mean square error (MSE), which is sometimes also known as the Brier score [47], has further been used in the imbalanced classification context. It is given by the usual formula:
@@
where, in the present context, $\widehat{p}_{i}$ is an estimated probability for a package’s release history having been affected by at least one vulnerability, while $p_{i}$ is the actual probability; $p_{i}=1$ in case the $i$:th package has been affected by at least one vulnerability. Unlike with BA and G-mean, lower values are better.
3.2.3 Computation
@@
The first model contains only releases. As was discussed in Subsection 3.1, it is the only explanatory metric that is not explicitly related to popularity of open source software. Therefore, the first model serves as a baseline for comparing the other two models. If the hypothesis holds, adding the popularity metrics should increase performance.
The second model contains two explanatory metrics: again releases and a sum variable based on the arithmetic mean of the six popularity metrics. Cronbach’s standardized $\alpha$-coefficient [49] is $0.64$ for this sum variable. Given the maximum of one, the value is not particularly high but still sufficient for the present purposes. Adjectives such as acceptable, satisfactory, and sufficient have been used to describe comparable values in existing research [50]. The third full or unrestricted model contains all available information in the dataset. In other words, the seven metrics described earlier in Subsection 3.1 are used in this full model. Alongside the hypothesis, the prior expectation is that the model’s performance is better compared to the barebone first model.
@@ 4 Results
4.1 Descriptive Statistics
The presentation of the results can be started by reiterating the methodological points. Thus, the extreme imbalance is best illustrated with plain numbers: of the over $381$ thousand PHP packages, only $777$, or about $0.2\%$, have been affected by one or more vulnerabilities over their release histories. When taking a look at Table 1, it can be further concluded that the majority (57%) of the $777$ packages with vulnerabilities has been affected by a single vulnerability. As has also been previously observed with PHP packages [12], the tail is relatively long; there are outlying packages that have been affected by nine or more vulnerabilities. A similar observation applies to the remaining metrics; they all have long tails. Thus, the $\ln(x+1)$ transformation is generally justified for analyzing the dataset.
Table 1: Vulnerabilities Across the PHP Packages

### v2

@@ Abstract
There has been a long-standing hypothesis that a software’s popularity is related to its security or insecurity in both research and popular discourse. There are also a few empirical studies that have examined the hypothesis, either explicitly or implicitly. The present work continues with and contributes to this research with a replication-motivated large-scale analysis of software written in the PHP programming language. The dataset examined contains nearly four hundred thousand open source software packages written in PHP. According to the results based on reported security vulnerabilities, the hypothesis holds; packages having seen reported vulnerabilities over their release histories are generally more popular than packages for which no vulnerabilities have been reported. With this replication results, the paper contributes to the efforts to strengthen the empirical knowledge basis in cyber and software security.
Index Terms:
@@
Software security, vulnerability, replication, software ecosystems, dependencies, downloads, installs, stars, forks
I Introduction
@@
The paper examines what can be called a popularity hypothesis in empirical software security research; an assumption that a software’s popularity affects its security or insecurity. The hypothesis has been previously examined with Java packages, the conclusion being that popularity is not a reliable indicator of software security [1]. Also another recent paper has evaluated the hypothesis implicitly, concluding that reported vulnerability counts do not correlate strongly with popularity metrics for Java packages [2]. The present paper revisits the hypothesis with a large-scale analysis of open source software packages written in the PHP programming language. Given the existing work, the paper is by definition a replication.
@@
Within the domain of replication studies, the paper can be classified as a dependent or a conceptual replication: although both the data and methods are different, the paper is designed with an explicit reference to previous work in mind [3, 4, 5, 6], including particularly the conclusions reached [7, 8]. The methodological difference can be justified on the grounds that the studies replicated were both based only on correlation analysis [1, 2]. In contrast, the paper uses also regression analysis and classification algorithms, which can be argued to yield stronger empirical evidence in general. Furthermore, the recently discussed generalizability problems in empirical software engineering [9] justify the use of a different dataset. If the hypothesis holds, or does not hold, with two distinct datasets and different methodologies, there is either stronger or weaker evidence for it, respectively.
@@
The underlying replication logic focuses on the previous work’s conclusion [7]—it is important to know whether or not the hypothesis holds with a different dataset and different methods because the hypothesis is rather fundamental in many ways. It also frequently appears in popular discourse. The primary historical example would be Microsoft Windows; not so long ago, it was commonly discussed and speculated by both laypeople and professionals that the popularity of the operating system was a definite factor behind its historical insecurity. To quote from a famous security engineering textbook, most “malware writers targeted Windows rather than Mac or Linux through the 2000s and 2010s as there are simply more Windows machines to infect”, and, furthermore, the “model replicated itself when smartphones took over the world” [10, p. 296] due to popularity of Android. Although there is no particular reason to question these assertions, it is still worth remarking that they are not actually backed by any empirical evidence in the textbook. This point motivates a need to examine the hypothesis further empirically.
@@
Because the two studies replicated did not support the hypothesis, the classical motivation for replication studies to try to falsify a theory [3], or a smaller hypothesis, takes the form of trying to empirically affirm that the hypothesis does in fact hold. Given a distinction between “hard” and “soft” theories in empirical software engineering [11], the popularity hypothesis is also on the side of hard theories because it presumes a rather direct relation between a software characteristic (popularity) and an effect (insecurity). However, as noted in the final Subsection V-D, the relation should not be interpreted as a strict causal relationship. This point is reinforced by emphasizing that the theorizing involved is still rather soft. In other words, assuming that the hypothesis holds, the underlying explanations remain open to a debate.
@@
The starting point for an explanation given by what is now known as (information or cyber) security economics would be incentives [10]. That is, cyber criminals and other adversaries have an incentive to target popular software products, packages, and projects, whether the targeting is about finding vulnerabilities from those or explicitly attacking them. Then: even though both commercial software vendors and open source software projects have an incentive to develop secure code and provide defenses, it is frequently discussed in the literature that software security and cyber security in general are not prioritized because there are no pressing incentives to do so [12]. A slightly different theorization path involves the famous so-called Linus law; that all bugs are shallow when there are many eyes involved to look at code [13]. From this alternative perspective, there are also empirical studies indicating that widely used and hence popular open source software packages often contain more reported bugs and vulnerabilities than less popular packages [14, 15, 16].
These empirical studies further reinforce the need and motivation for a replication of the popularity hypothesis studies already mentioned earlier. The italics placed upon the word reported are also worth briefly explaining. The reason for the italics is that empirical observations cannot generally prove that something is secure, although these can be used to make claims about insecurity [17]. This fundamental point should be kept in mind throughout the paper; only reported security issues are observed. By implication, the popularity hypothesis in itself may not necessarily tell everything, or in some cases even anything, about software security per se.
@@
    With respect to one of the replicated studies [1], the paper examines a substantially larger dataset. With respect to both replicated studies [1, 2], the paper uses more sophisticated methods for analyzing the large dataset.
@@
The paper’s remainder is structured into a conventional format. The opening Section II further frames the paper with a larger scope of related work. Then, the empirical materials and methods are elaborated in Section III. This section also clarifies what is meant by software popularity; in general, it is operationalized as a degree to which a package is widely adopted and actively engaged within its ecosystem. The subsequent Section IV presents the replication results after which conclusions follow in the final Section V.
II Related Work
The paper’s reference domain is empirical software engineering focusing on software security. The domain is large, but with some suitable framings, it is possible to narrow it.
@@ II Related Work
The first, obvious framing is toward empirical research. The second equally obvious framing is toward the PHP programming language. With these two framings, a few examples can be given about implicitly related work. For instance, technical debt of PHP packages for web applications has been investigated empirically [18]. Regarding web applications and websites in general, another example would be a large-scale empirical study investigating the adoption of PHP interpreter releases among website deployments [19]. A third framing can be done toward software security. The previous two examples are in the scope also with this further software security framing—technical debt is related to software quality, which is a superset for software security, while the adoption (or a lack thereof) of new PHP interpreter releases contains security risks in case many (or some important) websites do not update their interpreters for a reason or another. The reason for the security risks is that also the PHP interpreter itself has been affected many vulnerabilities over the years.
@@
A lot of empirical research has also been done to examine the security of software written in PHP more generally. For instance, the security of plugins for popular PHP-based web frameworks has been investigated [20, 16]. Regarding software vulnerabilities and their underlying weaknesses, and despite the availability of defensive solutions [21], cross-site scripting and input validation more generally have been the most typical weaknesses in PHP software [16, 22]. Though, it can be noted that this observation is hardly unique to PHP. In other words, the same weaknesses typically lead the scoreboards also with other popular interpreted programming languages used in the web domain [23]. Having said that, it should be emphasized that empirical software security research on PHP applications is not limited only on vulnerabilities and weaknesses. A good example would be a forensics investigation of PHP applications in containerized cloud environments based on empirical log analysis [24]. Furthermore, a fourth framing can be done toward replication research.
@@
A fifth and final clarifying framing can be done toward research on software ecosystems and particularly the security or insecurity of software distributed via them. The empirical security-oriented ecosystem research is again vast. A recurring theme is that software packages distributed in software ecosystems via language-specific package managers are generally of poor equality, often containing security issues of various kinds, including concrete vulnerabilities [27]. In addition, the ecosystems exhibit a risk of malware being uploaded, as is often done together with so-called typo-squatting [23, 28, 29]. The underlying security risks are intensified by the heavy use of software dependencies in these ecosystems [30]. Given this background, it is understandable that there are also various ongoing funded projects to improve the ecosystems and the supply chain security of open source software in general [12]. The software ecosystem research and dependencies are also important to underline because the concept of (software) popularity is closely related to them [31, 32]. The framing toward software ecosystems is also relevant as it is also directly related to the dataset examined. On that note, the materials and methods should also be elaborated.
III Materials and Methods
@@
III-A Data
The dataset was collected on December 2024 from Packagist [33], a repository for PHP software packaged with the Composer package manager [34]. The dataset is available online for any interested party, possibly regarding further replication work [35]. In total, as many as $n=381,993$ packages are covered in the dataset.
@@ III-A Data
The dataset is composed of eight simple meta-data metrics provided on the Packagist’s website: security, installs, dependents, suggesters, stars, watchers, forks, and releases. Of these metrics, all except security and releases can be interpreted to proxy slightly different dimensions of software popularity. The outlying releases metric counts the number of releases made for a given open source PHP package, as reported on the Packagist’s website. All metrics have a continuous scale.
@@
While installation amounts, installs, provide a relatively straightforward and widely used metric for software popularity [15, 36, 16, 31, 32], the dependents metric conveys a package’s popularity in terms of incoming edges in a dependency network between packages [31]. With this metric a package is seen as popular when many other packages depend on it. Both installation (or download) counts and dependencies have also been used to help security-related decision-making about funding applications for open source software projects [12]. Although documentation is unfortunately lacking, the suggesters metric generally proxies “recommended” dependencies—for each package, Packagist recommends also further packages in addition to a package’s conventional dependencies, as captured by dependents, without which the package does not install or function properly.
@@
The stars, watchers, and forks metrics are explicitly tied to functionality provided by GitHub on which almost all of the actual development of the PHP packages occurs today. In general, with “stars” developers can flag packages and projects they perceive as interesting or relevant. By adding oneself as a “watcher” to a project, one can get information about changes in the project. The forks metric, in turn, refers to the number of times a given project was forked, as has been common in the open source world throughout the decades [37]. According to practitioner surveys, all three metrics have been seen as relevant for proxying popularity [38]. Finally, the dependent metric is security, which counts and records vulnerabilities reported for a package. Many (but not all) of the reported vulnerabilities tracked on Packagist are identified with Common Exposures and Vulnerabilities (CVEs), while the associated security advisories typically point toward GitHub. In this regard, it can be noted that no additional data was retrieved to validate the reported CVE-referenced vulnerability information provided on Packagist. Nor was GitHub queried to retrieve further software development data.
III-B Methods
@@
The primary methodology is based on binary-valued classification for which security is truncated into a variable categorizing packages without any reported vulnerabilities and packages with one or more reported vulnerabilities in their whole release histories. In addition, a few descriptive statistics and a brief ordinary least squares (OLS) regression analysis are presented. As these do not require a particular exposition, the following discussion concentrates on the classification.
III-B1 Classification
@@
Imbalance is a highly typical and difficult problem in computational cyber security applications [39, 40, 41, 42]. Although not a problem for the current paper, not only is cyber security data often imbalanced but it also often contains missing values [43]. Even without a problem of missing values, as will be soon seen in Section IV, the dataset examined can be characterized to contain extreme imbalance. In other words, only a tiny minority of the PHP packages observed has been affected by a reported vulnerability or multiple reported vulnerabilities.
@@
To tackle the issue, four computational solutions are used. The first is oversampling or upsampling, as it is also known. It balances a training set by replicating the minority class, which in the present context refers to packages that have been affected by one or more reported vulnerabilities during their entire release histories. The second is undersampling or downsampling; it eliminates observations from the majority class (packages without no reported vulnerabilities) until a training set is balanced. Although there are variants that use synthetic data [40], both the oversampling and undersampling solutions are computed with random samples.
@@
III-B2 Performance Metrics
Imbalance affects also metrics for evaluating classification performance. Among other things [47], the traditional accuracy metric is biased and misleading in this context [53]. Although the $F_{1}$ metric is often used as an alternative, it too has been seen as problematic in the cyber security context [54]. Therefore, three alternative but still relatively well-known performance metrics are used.
@@ III-B2 Performance Metrics
Third, the conventional mean squared error (MSE), which is sometimes also known as the Brier score [53, 56], has further been used in the imbalanced classification context. It is given by the usual formula:
@@
where, in the present context, $\widehat{p}_{i}$ is an estimated probability for a package’s release history having seen at least one reported vulnerability, while $p_{i}$ is the actual probability; $p_{i}=1$ in case at least one vulnerability has been reported for the $i$:th package. Unlike with BA and G-mean, lower values are better.
III-B3 Computation
@@
The first model contains only releases. As was discussed in Subsection III-A, it is the only explanatory metric that is not explicitly related to popularity of open source software. Therefore, the first model serves as a baseline for comparing the other two models. If the hypothesis holds, adding the popularity metrics should increase performance.
The second model contains two explanatory metrics: again releases and a sum variable based on the arithmetic mean of the six popularity metrics. Cronbach’s standardized $\alpha$-coefficient [58] is $0.64$ for this sum variable. Given the maximum of one, the value is not particularly high but still sufficient for the present purposes. Adjectives such as acceptable, satisfactory, and sufficient have been used to describe comparable values in existing research [59]. The third full or unrestricted model contains all available information in the dataset. In other words, the seven metrics described earlier in Subsection III-A are used in this full model. Alongside the hypothesis, the prior expectation is that the model’s performance is better compared to the barebone first model.
@@ IV Results
IV-A Descriptive Statistics
The presentation of the results can be started by reiterating the methodological points. Thus, the extreme imbalance is best illustrated with plain numbers: of the over $381$ thousand PHP packages, only $777$, or about $0.2\%$, have seen one or more reported vulnerabilities over their release histories. When taking a look at Table I, it can be further concluded that a single vulnerability has been reported for the majority (57%) of the $777$ packages with a record of reported vulnerabilities. As has also been previously observed with PHP packages [16], however, the tail is relatively long; there are outlying packages that have seen nine or more reported vulnerabilities. A similar observation applies to the remaining metrics; they all have long tails. As can be seen from Fig. 1, also the popularity sum variable and releases both exhibit a long-tailed distribution too. This observation is familiar from existing studies [2, 23]. Thus, also the $\ln(x+1)$ transformation is generally justified.
TABLE I: Reported Vulnerabilities Across the PHP Packages